# Lossy Compression with Data, Perception, and Classification Constraints

### v2

@@ V-A DL-based Lossy Compression with RDPC Constraints
By utilizing a generative adversarial network (GAN), the process of loss compression $p_{\hat{X}|X}(\hat{x}|x)$ can be achieved with a encoder and a decoder parameterized by adjustable parameter $\phi_{e}$ and $\phi_{d}$ respectively. Let $\phi=\{\phi_{e},\phi_{d}\}$ denote the parameters of the whole encoder-decoder network.
@@
The reconstruction $\hat{x}$ and the original image $x$ are fed to the discriminator $G$ to compute the discriminator loss. The reconstruction $\hat{x}$ is also fed to a pre-trained classifier to obtain a one-hot vector $\hat{\mathbf{s}}$ with each entry being the probability that the image belongs to each class. The true label $s$ and the predicted $\mathbf{\hat{s}}$ are used to compute the classification accuracy.
@@
where the inequality follows from the non-negativity of KL- divergence, and $\text{CE}(s,\mathbf{\hat{s}})$ is the cross entropy loss.
@@
where $\lambda_{d},\lambda_{p},$ and $\lambda_{c}$ are hyperparameters to control weights of distortion, perception, and cross entropy losses, respectively.
@@ V-C Vanishment of Tradeoffs When Rate Increases
From Fig. 13 and Fig. 14, we observe that as the rate increases, both the $(D,C)$ and $(P,C)$ points become more concentrated. By setting a sufficiently high rate (e.g., $R=12544$ bits), the tradeoffs in RDC and RPC given D are almost gone. In this scenario, the classification accuracies consistently remain at 0.975, regardless of the MSE ranging from 0.006 to 0.014, and the W-1 loss remains consistently below 0.5.
Figure 13: The vanishment of RDC tradeoff given P when the rate $R$ increases for MNIST dataset.
Figure 14: The vanishment of RPC tradeoff given D when the rate $R$ increase for MNIST dataset.
VI Conclusions
@@ Appendix A Proof of Theorem 1
Consider the RDC function in (2) with Hamming distance for a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$).
@@
Note that by data-processing inequality[17, Theorem 2.8.1] and $S-X-\hat{X}$, we have
@@
We first find a lower bound of $I(X,\hat{X})$ leveraging the constraints and then show that it is achievable. By Mrs. Gerber’s Lemma proved in [18], when $H(S|\hat{X})\leq C$, we have $H(X|\hat{X})\leq H(C_{1})$, where $C_{1}\triangleq\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$ and $H^{-1}:[0,1]\rightarrow[0,\frac{1}{2}]$ denotes the inverse function of Shannon entropy for probability less than $\frac{1}{2}$. Note that $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}\leq\frac{\frac{1}{2}-p_{1}}{1-2p_{1}}=\frac{1}{2}$. Therefore, we have
@@ Appendix B Proof of Theorem 2
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[17, Chapter 8], the RDC problem can be represented by
@@
The second observation is that when $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the classification constraint is infeasible. To make $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ in (7) meaningful, we have $01-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}}>$ 0, i.e., $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}<\sigma_{x}^{2}$. Then, the mutual information between $S$ and $\hat{X}$ is upper bounded by
@@
making the (7b) infeasible if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$. Thus, in the following discussion, we assume $C\geq\frac{1}{2}\log(1-\frac{\theta_{1}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$.
In the case of (7b) not active, by Shannon’s classic rate-distortion theory for a Gaussian source with squared-error distortion[17, Chapter 10], the optimal rate is $\frac{1}{2}\log\frac{\sigma_{s}^{2}}{D}$ for $D\leq\sigma_{x}^{2}$, achieved by setting $\sigma_{\hat{x}}^{2}=\sigma_{x}^{2}-D$ and $\theta_{2}=\sigma_{x}^{2}-D$. For the classification constraint (7b), it is inactive if
@@
$R(D,C)=\begin{cases}\frac{1}{2}\log\frac{\sigma_{x}^{2}}{D}\\ \ \ \ \ \ \text{ for }D\leq\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ -\frac{1}{2}\log(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C}))\\ \ \ \ \ \ \text{ for }D>\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ 0\ \ \ \text{ for }C>h(S)\quad\text{and}\quad D>\sigma_{x}^{2},\end{cases}$
@@ Appendix C Proof of Theorem 3
Recall the RPC function in (3) for a Bernoulli source $X$ and a classification variable $S$ with binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$). Similar to the RDC case, if $C<H(S_{1})$, by data-processing inequality[17, Theorem 2.8.1], the problem is infeasible. In the following discussion, we assume $C\geq H(S_{1})$.
@@
Achievability: To achieve the optimal rate, we need to find the optimal solution $p_{a}=P(\hat{X}=0|X=0)$ and $p_{b}=P(\hat{X}=0|X=1)$. From the BSC assumption, we can obtained the marginal distribution of $X$ is $\text{Pr}(X=1)=\frac{a-p_{1}}{1-2p_{1}}\triangleq b$.
@@
At the same time, $H(S|\hat{X})=f(p_{a},\frac{(1-b)(1-p_{a})}{b})=g(p_{a})=C$ for the choice of $p_{a}=g^{-1}(C)$.
@@ Appendix D Proof of Theorem 4
The KL-divergence of two Gaussian varibales $X\sim\mathcal{N}(\mu_{x},\sigma^{2}_{x})$ and $\hat{X}\sim\mathcal{N}(\mu_{\hat{x}},\sigma^{2}_{\hat{x}})$ is $d_{\text{KL}}(p_{X},p_{\hat{X}})=\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}},$ where we take the natural base of logarithm.
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[17, Chapter 8], as well as the KL divergence of two Gaussian varibales, the RPC problem can be represented by
@@
Converse: Similar to the RDC case, by (10b) we have $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}\geq\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$, which gives us a lower bound of the objective function $I(X;\hat{X})-\frac{1}{2}\log\Big{(}1-\frac{\sigma_{s}^{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\Big{)}$, since $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ is an increasing function of $\frac{\theta_{2}^{2}}{\sigma^{2}_{\hat{x}}}$. The lower bound is achieved when $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$.
@@
Since $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$, we have $I(X;\hat{X})=-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})=-\frac{1}{2}\log\big{(}1-\frac{\sigma_{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\big{)}$ achieving the lower bound. Meanwhile, we have $\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}}=0<P$ since $\sigma_{\hat{x}}=\sigma_{x}$ and $\mu_{\hat{x}}=\mu_{x}$.

### v3

@@ V-A DL-based Lossy Compression with RDPC Constraints
By utilizing a generative adversarial network (GAN), the process of lossy compression $p_{\hat{X}|X}(\hat{x}|x)$ can be achieved with a encoder and a decoder parameterized by adjustable parameter $\phi_{e}$ and $\phi_{d}$ respectively. Let $\phi=\{\phi_{e},\phi_{d}\}$ denote the parameters of the whole encoder-decoder network.
@@
The reconstruction $\hat{x}$ and the original image $x$ are fed to the discriminator $G$ to compute the discriminator loss. The reconstruction $\hat{x}$ is also fed to a pre-trained classifier to obtain a vector $\hat{\mathbf{s}}$ with each entry being the probability that the image belongs to each class. The true label $s$ and the predicted $\mathbf{\hat{s}}$ are used to compute the classification accuracy.
@@
where the inequality follows from the non-negativity of KL- divergence, and $\text{CE}(s,\mathbf{\hat{s}})$ is the conditional cross-entropy loss.
Remark 5
In practical loss implementations, such as in PyTorch, people are using the same conditional cross-entropy loss as detailed above. Recall that $\hat{\mathbf{s}}$ is a vector output from the pre-trained classifier given $\hat{x}$, with the $i$ entry being the conditional probability of $i$th class, i.e., $p_{\psi}(s=i|\hat{x})$. Given parameters $\phi$, $\hat{x}=G_{\phi}(E_{\phi}(x))$ is a function of $x$. Denote the label of $x$ by $i_{x}$ and the batch size by $N$. Then $p_{\phi}(s,\hat{x})=p_{\phi}(s|G_{\phi}(E_{\phi}(x)))p_{\phi}(\hat{x})=1/N$ if $s=i_{x}$, and $p_{\phi}(s,\hat{x})=0$ otherwise. Thus, the cross-entropy loss can be expressed as
$\text{CE}(s,\mathbf{\hat{s}})$ $=\sum_{s}\!\sum_{\hat{x}}p_{\phi}(s,\hat{x})\!\log\!\frac{1}{p_{\psi}(s|\hat{x})}=-\frac{1}{N}\sum_{\hat{x}}\log p_{\psi}(i_{x}|\hat{x}),$
which is the implementation used in PyTorch.
@@
where $\lambda_{d},\lambda_{p},$ and $\lambda_{c}$ are hyperparameters to control weights of distortion, perception, and cross-entropy losses, respectively.
@@ V-C Vanishment of Tradeoffs When Rate Increases
From Fig. 13, we observe that as the rate increases, both the $(D,C)$ and $(P,C)$ points become more concentrated. By setting a sufficiently high rate (e.g., $R=12544$ bits), the tradeoffs in RDC and RPC given D are almost gone. In this scenario, the classification accuracies consistently remain at 0.975, regardless of the MSE ranging from 0.006 to 0.014, and the W-1 loss remains consistently below 0.5.
Figure 13: The vanishment of (a) RDC tradeoff given P and (b) RPC tradeoff given D when the rate $R$ increases for MNIST dataset.
VI Discussions
In previous sections, we characterized the properties of RDC and RPC functions for Binary and scalar Gaussian cases. We also explored the decisive role of distortion in the existence of tradeoffs. The above experiments empirically demonstrated the effectiveness of our theoretical results, verifying the characteristics of different tradeoffs on real-world datasets. In the following, we identify several concerns and suggest possible directions for future research to expand upon our framework.
VI-1 Joint consideration of rate, distortion, perception, and classification
In this paper, we consider two ternary tradeoffs, namely RDC and RPC. The joint consideration of four aspects of rate, distortion, perception, and classification remains unresolved. Note that the techniques involved in the derivation of closed-form RDC and RPC diverge from those of RDP and IB problems, which involve carefully determining the activeness boundary or delicately finding a solution with perfect perception. Consequently, extension to the case with all three constraints (1a)-(1c) is not straightforward, even in binary or scalar Gaussian scenarios. A primary attempt may involve studying the boundary at which certain constraints become inactive, which is more complicated than the ternary case. Recent studies have shown that the RDP function and IB problem can be formulated as an optimal transport problem [46, 47], and then the critical transition can be identified [39]. Analogous methodologies may be applied to the RDPC function. Meanwhile, techniques from network information theory [48] may also be used to characterize the bound of the RDPC function.
VI-2 Extension to other source distributions
The theoretical results of RDC and RPC functions mainly focus on Binary and scalar Gaussian cases, which is a strong foundation. However, it is plausible to derive iterative algorithms to characterize the RDC, RPC, and even RDPC functions for general source distributions. In [2], the authors proposed to solve the IB problem with a natural extension of the Blahut-Arimoto (BA) algorithm based on self-consistent equations. In [49, Section III-A], a modified version of the BA algorithm is also applied to compute the optimal solution of the RDP problem for general sources, suggesting the potential to extend such alternating minimization process to our RDC, RPC, or RDPC problems. However, such an extension is not straightforward since the convergence and correctness of the algorithms need to be further verified. Meanwhile, since RDP and IB have been formulated as optimal transport problems [46, 39] and can be solved by modified sinkhorn algorithms[47, 50], similar techniques could potentially be applied to tackle our RDC, RPC, and RDPC functions.
VI-3 Extension to general distortion and other noise models
In practice, many deep learning-based methods utilize more general distortion metrics beyond pixel-level MSE. For example, feature reconstruction loss [51] computes the feature-level MSE, where the features correspond to the outputs of the $j$th layer of a VGG-16 network. By incorporating more generalized distortion and perceptual metrics into the RDPC framework, we can broaden the scope of applications for our theories. Meanwhile, we re-examined the tradeoffs under additive Gaussian noise in Section IV-A. In forthcoming studies, we can explore more general degradation scenarios, such as fading channels, to make our theory more applicable to practical wireless communication systems.
VI-4 Traversing different tradeoffs in real time
In real-world communication systems, dynamically navigating tradeoffs in real time is vital to meet diverse task demands. In prior works [21, 9], the authors proposed to employ generative adversarial networks (GAN) to traverse the DP tradeoff and the RDP function. However, to achieve different tradeoffs between rate, distortion, and perception, one needs to adjust loss functions and retrain the model, which is resource-consuming and impractical in real-world systems. Recent advances in image denoising[23, 52] have made progress on traversing DP tradeoff at inference time for specific tasks, e.g., super-resolution or Gaussian deblurring. Nevertheless, the flexible navigation of DP tradeoffs for general inverse problems remains an open challenge, let alone for the RDP and newly proposed RDC and RPC tradeoffs.
VII Conclusions
@@ Appendix A Proof of Theorem 1
Consider the RDC function in (2) with Hamming distance for a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq 1/2$).
@@
Note that by data-processing inequality[18, Theorem 2.8.1] and $S-X-\hat{X}$, we have
@@
We first find a lower bound of $I(X,\hat{X})$ leveraging the constraints and then show that it is achievable. By Mrs. Gerber’s Lemma proved in [19], when $H(S|\hat{X})\leq C$, we have $H(X|\hat{X})\leq H(C_{1})$, where $C_{1}\triangleq\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$ and $H^{-1}:[0,1]\rightarrow[0,1/2]$ denotes the inverse function of Shannon entropy for probability less than $\frac{1}{2}$. Note that $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}\leq\frac{\frac{1}{2}-p_{1}}{1-2p_{1}}=\frac{1}{2}$. Therefore, we have
@@ Appendix B Proof of Theorem 2
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[18, Chapter 8], the RDC problem can be represented by
@@
The second observation is that when $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the classification constraint is infeasible. To make $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ in (7) meaningful, we have $01-\theta_{2}^{2}/({\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})>$ 0, i.e., $\theta_{2}^{2}/\sigma_{\hat{x}}^{2}<\sigma_{x}^{2}$. Then, the mutual information between $S$ and $\hat{X}$ is upper bounded by
@@
making the (7b) infeasible if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$. Thus, in the following discussion, we assume $C\geq\frac{1}{2}\log(1-\theta_{1}/({\sigma_{s}^{2}\sigma_{x}^{2}}))+h(S)$.
In the case of (7b) not active, by Shannon’s classic rate-distortion theory for a Gaussian source with squared-error distortion[18, Chapter 10], the optimal rate is $\frac{1}{2}\log\frac{\sigma_{s}^{2}}{D}$ for $D\leq\sigma_{x}^{2}$, achieved by setting $\sigma_{\hat{x}}^{2}=\sigma_{x}^{2}-D$ and $\theta_{2}=\sigma_{x}^{2}-D$. For the classification constraint (7b), it is inactive if
@@
$R(D,C)=\begin{cases}\frac{1}{2}\log\frac{\sigma_{x}^{2}}{D}~{}&\text{ for }D\leq\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ -\frac{1}{2}\log(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C}))~{}&\text{ for }D>\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ 0\ \ &\text{ for }C>h(S)\quad\text{and}\quad D>\sigma_{x}^{2},\end{cases}$
@@ Appendix C Proof of Theorem 3
Recall the RPC function in (3) for a Bernoulli source $X$ and a classification variable $S$ with binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$). Similar to the RDC case, if $C<H(S_{1})$, by data-processing inequality[18, Theorem 2.8.1], the problem is infeasible. In the following discussion, we assume $C\geq H(S_{1})$.
@@
Achievability: To achieve the optimal rate, we need to find the optimal solution $p_{a}=P(\hat{X}=0|X=0)$ and $p_{b}=P(\hat{X}=0|X=1)$. From the BSC assumption, we can obtained the marginal distribution of $X$ is $P(X=1)=(a-p_{1})/(1-2p_{1})\triangleq b$.
@@
At the same time, we have the classification constraint $H(S|\hat{X})=f(p_{a},(1-b)(1-p_{a})/b)\!=g(p_{a})=C$ for the choice of $p_{a}=g^{-1}(C)$.
@@ Appendix D Proof of Theorem 4
The KL-divergence of two Gaussian varibales $X\sim\mathcal{N}(\mu_{x},\sigma^{2}_{x})$ and $\hat{X}\sim\mathcal{N}(\mu_{\hat{x}},\sigma^{2}_{\hat{x}})$ is
$d_{\text{KL}}(p_{X},p_{\hat{X}})=\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}},$
where we take the natural base of logarithm.
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[18, Chapter 8], as well as the KL divergence of two Gaussian varibales, the RPC problem can be represented by
@@
Converse: Similar to the RDC case, by (10b) we have
$\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}\geq\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C}),$
which gives us a lower bound of the objective function:
$I(X;\hat{X})-\frac{1}{2}\log\big{(}1-\frac{\sigma_{s}^{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\big{)},$
since $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ is an increasing function of $\frac{\theta_{2}^{2}}{\sigma^{2}_{\hat{x}}}$. The lower bound is achieved when $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$.
@@
Since $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$, we have
$I(X;\hat{X})$ $=-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})=-\frac{1}{2}\log\big{(}1-\frac{\sigma_{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\big{)}$
achieving the lower bound. Meanwhile, we have $\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}}=0<P$ since $\sigma_{\hat{x}}=\sigma_{x}$ and $\mu_{\hat{x}}=\mu_{x}$.